Calculate the hydroxide Ion concentration of an aqueous solution having pH 4 .
1
Expert's answer
2019-11-18T06:29:07-0500

GIVEN

pH = 4


WE KNOW


pH + pOH = 14


pOH = 14 - 4


pOH = 10


WE ALSO KNOW THAT


pOH = -log [OH - ]


SO ,

-log10 [OH - ] = 10


[OH - ] = 10-10


HENCE ,

HYDROXIDE ION CONCENTRATION OF AN AQUEOUS SOLUTION HAVING PH 4 IS [OH - ] = 10 -10